Given 11, 6**** 1x2 + y2 dydx. Rewrite the integral in polar coordinates. Select one: radrde O b. Card ST S Srdr ST rdrde rP Please quickly and clear  
Answer #1

in polar coordinates x = rcos\small \theta

and y = rsin\small \theta

x2 + y2 = r2

\small \sqrt{x^{2}+y^{2}} =r denotes the upper half of semicircle

1593276785077_image.png

and for upper half of circle r varies from 0 to 1 and theta varies from 0 to pi

when varies from x = 1 to -1 theta varies from 0 to pi

so the double integration changes

\small \int_{0}^{\pi }\int_{0}^{1}r^{2} dr d\theta

option d is correct